OpenStudy (31356):

64-40ab=8(8-5ab)? MEDAL REWARDED!!

OpenStudy (31356):

@jim_thompson5910

OpenStudy (anonymous):

64-40ab= (8*8)-(8*5ab)

OpenStudy (31356):

Thanks Larry!

jimthompson5910 (jim_thompson5910):

you are correct again check: 64-40ab=8(8-5ab) 64-40ab=8(8)+8(-5ab) 64-40ab=64-40ab ... answer confirmed

OpenStudy (31356):

xy^2+xy=xy(y+1)

OpenStudy (31356):

Thanks jim

OpenStudy (anonymous):

64-40ab=64-40ab, 64=64

OpenStudy (31356):

Thanks, can you help me with the other one?

OpenStudy (anonymous):

xy^2+xy= xy(y) + xy(1)

OpenStudy (31356):

I think it is correct is it?

OpenStudy (anonymous):

xy(y) = xy^2; xy(1) = xy Yes

OpenStudy (31356):

Thanks Jim!

OpenStudy (anonymous):

Kind of odd though, the equation equals itself instead of you solving for a variable.

OpenStudy (31356):

Does it mean it's correct? Not so sure about this one, what do you think @jim_thompson5910 ?

OpenStudy (anonymous):

Both sides of the equation are equal to each other.

OpenStudy (31356):

Thanks!